# 360+x: A Panoptic Multi-modal Scene Understanding Dataset

### v1

@@ Abstract
Human perception of the world is shaped by a multitude of viewpoints and modalities. While many existing datasets focus on scene understanding from a certain perspective (e.g. egocentric or third-person views), our dataset offers a panoptic perspective (i.e. multiple viewpoints with multiple data modalities). Specifically, we encapsulate third-person panoramic and front views, as well as egocentric monocular/binocular views with rich modalities including video, multi-channel audio, directional binaural delay, location data and textual scene descriptions within each scene captured, presenting comprehensive observation of the world. Figure 1 offers a glimpse of all 28 scene categories of our 360+x dataset. To the best of our knowledge, this is the first database that covers multiple viewpoints with multiple data modalities to mimic how daily information is accessed in the real world. Through our benchmark analysis, we presented 5 different scene understanding tasks on the proposed 360+x dataset to evaluate the impact and benefit of each data modality and perspective in panoptic scene understanding. We hope this unique dataset could broaden the scope of comprehensive scene understanding and encourage the community to approach these problems from more diverse perspectives.
@@
1.1[Uncaptioned image]
Figure 1: Example 360panoramics videos from all 28 scene categories.
@@ 1 Introduction
Taking the above observations into consideration, a new dataset covering all these aforementioned aspects is presented in this work, to provide a panoptic scene understanding, termed 360+x dataset. This new dataset offers a diverse selection of perspectives, including a 360 panoramic view providing a complete panoptic view of the environment, and a third-person front view that highlights the region of interest that has the most movements in front of the camera. Additionally, we have included egocentric monocular and binocular videos to capture the first-person perspective of individuals in the environment. These viewpoints are complemented by aligned multi-channel audio with directional binaural delay information, as well as location information and scene descriptions as metadata. An illustration of the presented dataset collection system is shown in Figure 2.
Figure 2: Illustration of the proposed 360+x dataset. The 360 camera records fish-eye raw videos with front and back lenses. These videos are merged to create a spherical 360panorama (middle-up figure, zoom in for details), which is then transformed to (a) 360panoramic data using equirectangular projection. The (b) third-person front view is obtained by de-warping the rich movements region highlighted red in the spherical field of 360 panorama (the middle-left figure). By wearing stereo cameras, the capturers record (c) egocentric clips while staying visible to the fixed 360camera (central ellipse). (e) Directional audio time delay data is generated from left and right audio inputs (d) from the 360camera by interaural time delay process [4]. This helps locate sound sources in the 360 panorama.
@@ 3.1 Data Acquisition and Alignment
Egocentric binocular videos, as shown in Figure 2(c), were captured ranging from approximately 30 seconds to 1 minute in duration for each clip. A total of 1 to 5 stereo clips were recorded, scattered throughout the duration of the average 6 mins 360 video. In addition to stereo videos, we also provide the corresponding monocular videos for the egocentric view.
The audio recordings were temporally aligned with their corresponding videos with left/right channel modality. The four-channel audios with the 360 panoramic video are provided as well for further exploration. Moreover, we also provide the directional information of the audio which was presented using the estimated interaural time delay of the sound obtained from the method introduced in [4]. The GPS information and weather information were also provided.
@@ 3.2 Scene Selection
Third-person 360 Ego Ego Normal Directional GPS Avg Total Frames Annotations Multiple
@@ 3.4 Dataset Statistics and Analysis
Existing publicly available datasets primarily focus on visual unimodality [28, 16, 18, 7, 6]. In contrast, our dataset introduces a novel approach by collecting different views or modalities, as presented in Table 1, including 360 panoramic video, third-person front view video, egocentric monocular video, egocentric binocular video, normal audio, directional binaural delay, location and textual scene description. This diverse range of modalities provides multiple dimensions and clues for understanding and analysing complex scenes. Our dataset consists of 2,152 videos representing 232 data examples, with 464 videos captured using the 360 camera and the remaining 1,688 recorded with the Spectacles camera.
@@ 4 Benchmark and Experiments
Remark: Unless specifically stated otherwise, the experiments on 360+x will utilise three views: the 360 view, egocentric binocular view, and the third-person front view.
@@ 4.1 Experimental Setting
Table 2: Video classification performance across different views (Ego: egocentric binocular view, Front: third-person front view, and 360: 360 view) and data modalities (V: Video, A: Audio, D: Directional binaural delay). Reported in Avg. Prec. (%).
@@ V V + A V + A + D
360 Only 56.33  ($+$4.4) 67.14  ($+$11.9) 70.95  ($+$12.0)
360 + Egocentric 58.99  ($+$7.0) 70.48  ($+$15.2) 72.11  ($+$13.2)
360 + Front 59.70  ($+$7.8) 75.06  ($+$19.8) 77.69  ($+$18.8)
360 + Front + Ego 63.73  ($+$11.8) 77.32  ($+$22.1) 80.62  ($+$21.7)
@@ 4.2 Video Scene Classification
First, we are interested in the influence of different combinations of video views on the classification performance. The results, representing each combination, are summarised in Table 2. The results for single views are presented in the first three rows, indicating that using a single 360 panoramic view outperforms using either an egocentric binocular view or a third-person front view only. When employing multiple views, it is noted that better performance can be achieved compared to using a single view. Specifically, utilising all three views leads to the best performance. Such a performance can be attributed to the fact that although these three views describe the same scene, each different view offers a unique perspective that contributes to a more comprehensive understanding of the scene, resulting in improved performance.
@@ 4.3 Temporal Action Localisation
We provide a concise overview of the performance comparison for various temporal action localisation methods, including ActionFormer [40], TriDet [26] and TemporalMaxer [29], between the baseline extractors and our Pre-trained on 360+x extractors. The summarised results are presented in Table 3, from which we can see that the introduction of additional modalities (i.e. audio and direction binaural delay) has a prominent positive impact on the TAL task, leading to performance improvements for both sets of extractors. This result highlights the importance of leveraging multiple modalities in enhancing the accuracy and effectiveness of temporal activity localisation techniques. Using our custom extractors can provide additional improvements, as the baseline extractors may not be optimised for our specific binocular or 360 views. Additional results on variations of views can be found in the supplementary material section G.1.
@@ 4.4 Cross-modality Retrieval
Table 5 Models with different pre-train methods were fine-tuned and tested on video classification. The experiments use all three video views. Reported in Avg. Prec. (%).
@@ V V + A V + A + D
Table 6 Comparison between supervised pre-trained extractors with SSL pre-trained counterparts on TAL task. The experiments use all three video views with modalities (V+A+D).
@@ 4.5 Self-supervised Representation Learning
We first examined the impact of self-supervised learning models for video classification. Table 5  demonstrates the consistent precision gains achieved by utilising SSL pre-trained models. Notably, leveraging both video pace and clip order SSL techniques resulted in an average performance improvement of $\sim$ 7%.
We proceeded to perform experiments using SSL pre-trained models as feature extractors for the temporal action localisation task incorporating all three modalities (V+A+D) with the TriDet framework [26]. Since a training-from-scratch model cannot serve as the first-stage extractor, we employed the supervised extractors from section 4.2 as a comparison. The summarised results in Table 6  indicate that pre-training with video pace (VP) or clip order (CO) individually leads to an average performance improvement of $\sim$ 1.2% and $\sim$ 0.9% respectively on average, compared to the supervised baseline. The combination of both SSL methods yields the highest performance gain of $\sim$ 1.9%.
@@ Appendix E Explain of Self-supervised Learning
It is noteworthy that the VP and CO primarily focus on leveraging temporal information as training guidance, applying it either globally (pace) or locally (clip) to offer distinct interventions to this temporal data. By combining these interventions, there is potential to enhance the model’s capability to capture global and local temporal dependencies simultaneously. This integration, depicted in Figure 5, is delineated as ‘combine clip order and pace prediction’ or varied pace clip order (VP+PO) shuffle. This integration is highlighted in our experiments detailed in the main paper Tables 5  and 6 , where noted benefits become evident.
@@ G.1 Temporal Action Localisation
360Only $21.1$ $15.3$ $5.5$ $14.0$ ($+$5.1) $26.4$ $18.5$ $6.9$ $17.3$ ($+$6.3) $27.1$ $18.7$ $7.0$ $17.6$ ($+$6.2)
360+ Egocentric $21.4$ $15.8$ $5.7$ $14.3$ ($+$5.4) $27.3$ $19.2$ $7.2$ $17.9$ ($+$6.9) $27.8$ $19.6$ $7.2$ $18.2$ ($+$6.8)
360+ Front $24.2$ $16.8$ $6.1$ $15.7$ ($+$6.8) $28.1$ $20.3$ $7.3$ $18.6$ ($+$7.6) $28.2$ $20.8$ $7.3$ $18.8$ ($+$7.4)
360+ Front + Ego $24.6$ $17.1$ $6.3$ $16.0$  ($+$7.1) $28.2$ $20.6$ $7.3$ $18.7$  ($+$7.7) $28.8$ $21.0$ $7.4$ $19.1$  ($+$7.7)
@@ G.5 Transformer-Based Backbone
360Only , with I3D $16.7$ ($\pm$0.0) $10.1$ ($\pm$0.0) $4.8$ ($\pm$0.0) $10.5$ ($\pm$0.0) $23.6$ ($\pm$0.0) $17.2$ ($\pm$0.0) $6.4$ ($\pm$0.0) $15.7$ ($\pm$0.0)
360Only $17.1$ ($+$$0.4$) $13.4$ ($+$$3.3$) $5.2$ ($+$$0.4$) $11.9$ ($+$$1.4$) $25.9$ ($+$$2.3$) $18.5$ ($+$$1.3$) $6.1$ (--$0.3$) $16.8$ ($+$$1.1$)
360+ Egocentric $16.9$ ($+$0.2) $13.1$ ($+$3.0) $5.0$ ($+$0.2) $11.7$ ($+$1.1) $26.4$ ($+$2.8) $19.0$ ($+$1.8) $6.2$ (--0.2) $17.2$ ($+$1.5)
360+ Front $19.5$ ($+$2.8) $16.3$ ($+$6.2) $5.6$ ($+$0.8) $13.8$ ($+$3.3) $27.6$ ($+$4.0) $21.2$ ($+$4.0) $6.5$ ($+$0.1) $18.4$ ($+$2.7)
360+ Front + Ego $19.2$  ($+$2.5) $15.8$  ($+$5.7) $5.4$  ($+$0.6) $13.5$  ($+$2.9) $27.8$  ($+$4.2) $21.7$  ($+$4.5) $6.6$  ($+$0.2) $18.7$  ($+$3.0)
@@ Appendix I Social Impact
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example
@@
(a) 360panoramic video frame example

### v2

@@ Abstract
Human perception of the world is shaped by a multitude of viewpoints and modalities. While many existing datasets focus on scene understanding from a certain perspective (e.g. egocentric or third-person views), our dataset offers a panoptic perspective (i.e. multiple viewpoints with multiple data modalities). Specifically, we encapsulate third-person panoramic and front views, as well as egocentric monocular/binocular views with rich modalities including video, multi-channel audio, directional binaural delay, location data and textual scene descriptions within each scene captured, presenting comprehensive observation of the world. Figure 360+x: A Panoptic Multi-modal Scene Understanding Dataset offers a glimpse of all 28 scene categories of our 360+x dataset. To the best of our knowledge, this is the first database that covers multiple viewpoints with multiple data modalities to mimic how daily information is accessed in the real world. Through our benchmark analysis, we presented 5 different scene understanding tasks on the proposed 360+x dataset to evaluate the impact and benefit of each data modality and perspective in panoptic scene understanding. We hope this unique dataset could broaden the scope of comprehensive scene understanding and encourage the community to approach these problems from more diverse perspectives.
@@
1.1[Uncaptioned image] Figure 1: Example ${}^{\circ}$panoramics videos from all 28 scene categories.
@@ 1 Introduction
Taking the above observations into consideration, a new dataset covering all these aforementioned aspects is presented in this work, to provide a panoptic scene understanding, termed 360+x dataset. This new dataset offers a diverse selection of perspectives, including a ${}^{\circ}$ panoramic view providing a complete panoptic view of the environment, and a third-person front view that highlights the region of interest that has the most movements in front of the camera. Additionally, we have included egocentric monocular and binocular videos to capture the first-person perspective of individuals in the environment. These viewpoints are complemented by aligned multi-channel audio with directional binaural delay information, as well as location information and scene descriptions as metadata. An illustration of the presented dataset collection system is shown in Figure 2.
Figure 2: Illustration of the proposed 360+x dataset. The ${}^{\circ}$ camera records fish-eye raw videos with front and back lenses. These videos are merged to create a spherical ${}^{\circ}$panorama (middle-up figure, zoom in for details), which is then transformed to (a) ${}^{\circ}$panoramic data using equirectangular projection. The (b) third-person front view is obtained by de-warping the rich movements region highlighted red in the spherical field of ${}^{\circ}$ panorama (the middle-left figure). By wearing stereo cameras, the capturers record (c) egocentric clips while staying visible to the fixed ${}^{\circ}$camera (central ellipse). (e) Directional audio time delay data is generated from left and right audio inputs (d) from the ${}^{\circ}$camera by interaural time delay process [4]. This helps locate sound sources in the ${}^{\circ}$ panorama.
@@ 3.1 Data Acquisition and Alignment
Egocentric binocular videos, as shown in Figure 2(c), were captured ranging from approximately 30 seconds to 1 minute in duration for each clip. A total of 1 to 5 stereo clips were recorded, scattered throughout the duration of the average 6 mins ${}^{\circ}$ video. In addition to stereo videos, we also provide the corresponding monocular videos for the egocentric view.
The audio recordings were temporally aligned with their corresponding videos with left/right channel modality. The four-channel audios with the ${}^{\circ}$ panoramic video are provided as well for further exploration. Moreover, we also provide the directional information of the audio which was presented using the estimated interaural time delay of the sound obtained from the method introduced in [4]. The GPS information and weather information were also provided.
@@ 3.2 Scene Selection
Third-person ${}^{\circ}$ Ego Ego Normal Directional GPS Avg Total Frames Annotations Multiple
@@ 3.4 Dataset Statistics and Analysis
Existing publicly available datasets primarily focus on visual unimodality [28, 16, 18, 7, 6]. In contrast, our dataset introduces a novel approach by collecting different views or modalities, as presented in Table 1, including ${}^{\circ}$ panoramic video, third-person front view video, egocentric monocular video, egocentric binocular video, normal audio, directional binaural delay, location and textual scene description. This diverse range of modalities provides multiple dimensions and clues for understanding and analysing complex scenes. Our dataset consists of 2,152 videos representing 232 data examples, with 464 videos captured using the 360 camera and the remaining 1,688 recorded with the Spectacles camera.
@@ 4 Benchmark and Experiments
Remark: Unless specifically stated otherwise, the experiments on 360+x will utilise three views: the ${}^{\circ}$ view, egocentric binocular view, and the third-person front view.
@@ 4.1 Experimental Setting
Table 2: Video classification performance across different views (Ego: egocentric binocular view, Front: third-person front view, and ${}^{\circ}$: ${}^{\circ}$ view) and data modalities (V: Video, A: Audio, D: Directional binaural delay). Reported in Avg. Prec. (%).
@@ V V + A V + A + D
${}^{\circ}$ Only 56.33  ($+$4.4) 67.14  ($+$11.9) 70.95  ($+$12.0)
${}^{\circ}$ + Egocentric 58.99  ($+$7.0) 70.48  ($+$15.2) 72.11  ($+$13.2)
${}^{\circ}$ + Front 59.70  ($+$7.8) 75.06  ($+$19.8) 77.69  ($+$18.8)
${}^{\circ}$ + Front + Ego 63.73  ($+$11.8) 77.32  ($+$22.1) 80.62  ($+$21.7)
@@ 4.2 Video Scene Classification
First, we are interested in the influence of different combinations of video views on the classification performance. The results, representing each combination, are summarised in Table 2. The results for single views are presented in the first three rows, indicating that using a single ${}^{\circ}$ panoramic view outperforms using either an egocentric binocular view or a third-person front view only. When employing multiple views, it is noted that better performance can be achieved compared to using a single view. Specifically, utilising all three views leads to the best performance. Such a performance can be attributed to the fact that although these three views describe the same scene, each different view offers a unique perspective that contributes to a more comprehensive understanding of the scene, resulting in improved performance.
@@ 4.3 Temporal Action Localisation
We provide a concise overview of the performance comparison for various temporal action localisation methods, including ActionFormer [40], TriDet [26] and TemporalMaxer [29], between the baseline extractors and our Pre-trained on 360+x extractors. The summarised results are presented in Table 3, from which we can see that the introduction of additional modalities (i.e. audio and direction binaural delay) has a prominent positive impact on the TAL task, leading to performance improvements for both sets of extractors. This result highlights the importance of leveraging multiple modalities in enhancing the accuracy and effectiveness of temporal activity localisation techniques. Using our custom extractors can provide additional improvements, as the baseline extractors may not be optimised for our specific binocular or ${}^{\circ}$ views. Additional results on variations of views can be found in the supplementary material section G.1.
@@ 4.4 Cross-modality Retrieval
[…]
@@ V V + A V + A + D
[…]
@@ 4.5 Self-supervised Representation Learning
We first examined the impact of self-supervised learning models for video classification. Table 4.4  demonstrates the consistent precision gains achieved by utilising SSL pre-trained models. Notably, leveraging both video pace and clip order SSL techniques resulted in an average performance improvement of $\sim$ 7%.
We proceeded to perform experiments using SSL pre-trained models as feature extractors for the temporal action localisation task incorporating all three modalities (V+A+D) with the TriDet framework [26]. Since a training-from-scratch model cannot serve as the first-stage extractor, we employed the supervised extractors from section 4.2 as a comparison. The summarised results in Table 4.4  indicate that pre-training with video pace (VP) or clip order (CO) individually leads to an average performance improvement of $\sim$ 1.2% and $\sim$ 0.9% respectively on average, compared to the supervised baseline. The combination of both SSL methods yields the highest performance gain of $\sim$ 1.9%.
@@ Appendix E Explain of Self-supervised Learning
It is noteworthy that the VP and CO primarily focus on leveraging temporal information as training guidance, applying it either globally (pace) or locally (clip) to offer distinct interventions to this temporal data. By combining these interventions, there is potential to enhance the model’s capability to capture global and local temporal dependencies simultaneously. This integration, depicted in Figure 5, is delineated as ‘combine clip order and pace prediction’ or varied pace clip order (VP+PO) shuffle. This integration is highlighted in our experiments detailed in the main paper Tables 4.4  and 4.4 , where noted benefits become evident.
@@ G.1 Temporal Action Localisation
${}^{\circ}$nly $21.1$ $15.3$ $5.5$ $14.0$ ($+$5.1) $26.4$ $18.5$ $6.9$ $17.3$ ($+$6.3) $27.1$ $18.7$ $7.0$ $17.6$ ($+$6.2)
${}^{\circ}$+ Egocentric $21.4$ $15.8$ $5.7$ $14.3$ ($+$5.4) $27.3$ $19.2$ $7.2$ $17.9$ ($+$6.9) $27.8$ $19.6$ $7.2$ $18.2$ ($+$6.8)
${}^{\circ}$+ Front $24.2$ $16.8$ $6.1$ $15.7$ ($+$6.8) $28.1$ $20.3$ $7.3$ $18.6$ ($+$7.6) $28.2$ $20.8$ $7.3$ $18.8$ ($+$7.4)
${}^{\circ}$+ Front + Ego $24.6$ $17.1$ $6.3$ $16.0$  ($+$7.1) $28.2$ $20.6$ $7.3$ $18.7$  ($+$7.7) $28.8$ $21.0$ $7.4$ $19.1$  ($+$7.7)
@@ G.5 Transformer-Based Backbone
${}^{\circ}$nly , with I3D $16.7$ ($\pm$0.0) $10.1$ ($\pm$0.0) $4.8$ ($\pm$0.0) $10.5$ ($\pm$0.0) $23.6$ ($\pm$0.0) $17.2$ ($\pm$0.0) $6.4$ ($\pm$0.0) $15.7$ ($\pm$0.0)
${}^{\circ}$nly $17.1$ ($+$$0.4$) $13.4$ ($+$$3.3$) $5.2$ ($+$$0.4$) $11.9$ ($+$$1.4$) $25.9$ ($+$$2.3$) $18.5$ ($+$$1.3$) $6.1$ (--$0.3$) $16.8$ ($+$$1.1$)
${}^{\circ}$+ Egocentric $16.9$ ($+$0.2) $13.1$ ($+$3.0) $5.0$ ($+$0.2) $11.7$ ($+$1.1) $26.4$ ($+$2.8) $19.0$ ($+$1.8) $6.2$ (--0.2) $17.2$ ($+$1.5)
${}^{\circ}$+ Front $19.5$ ($+$2.8) $16.3$ ($+$6.2) $5.6$ ($+$0.8) $13.8$ ($+$3.3) $27.6$ ($+$4.0) $21.2$ ($+$4.0) $6.5$ ($+$0.1) $18.4$ ($+$2.7)
${}^{\circ}$+ Front + Ego $19.2$  ($+$2.5) $15.8$  ($+$5.7) $5.4$  ($+$0.6) $13.5$  ($+$2.9) $27.8$  ($+$4.2) $21.7$  ($+$4.5) $6.6$  ($+$0.2) $18.7$  ($+$3.0)
@@ Appendix I Social Impact
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example
@@
(a) ${}^{\circ}$panoramic video frame example